# 6G: The Intelligent Network of Everything – A Comprehensive Vision, Survey, and Tutorial

### v2

@@ VIII-C2 Advanced Modulation and Coding Methods
  • Past and Present: Each mobile generation has adopted a novel or enhanced channel coding concept, i.e., 2G: convolutional coding (data and control), 3G: Turbo coding (data) and convolutional coding (control), 4G: Turbo coding (data) and convolutional coding (control), 5G: low-density parity check (LDPC) coding (data), and polar coding (control) [352]. Currently, LDPC codes are used for data channels and polar codes for control channels in 5G [353]. Turbo coding was the first channel coding method approaching theoretical channel capacity. Linear LDPC coding also belongs to the class of capacity-approaching codes while outperforming Turbo coding, particularly in the higher coding rate region and in terms of the error floor [354]. Polar coding is a linear block-error correction code with low-complexity encoding and decoding designs, and the first channel code that achieves theoretical channel capacity [354]. LDPC and polar codes have been reviewed for 5G NR in [353, 354].
@@
    At a high level, there are two main challenges in modulation and coding designs toward 6G, i.e., developing efficient methods for ultra-high throughput and extremely reliable communication links. The former design problem calls for efficient high-order modulation and low coding rate methods with decent error-correcting properties. A straightforward solution is to increase the modulation order beyond 5G’s 1024-QAM, requiring a thorough performance comparison between different variations of QAM. However, increasing the order of QAM may not provide the needed performance gains at higher 6G frequencies due to the inefficiency of power amplifiers [63]. To alleviate this problem, coded modulation with signal shaping is a promising concept [358], especially in combination with polar codes [359, 360, 361, 362]. Polar codes have favorable features to efficiently support lower coding rates in combination with higher-order modulations [359, 363, 364]. Coded modulation with probabilistic shaping is already in use in optical fiber communications [63].
@@
  • Literature and Future Directions: In the 2010s, LDPC and polar codes were studied for 5G [353]. Research beyond 5G is ongoing, with a special emphasis on enhanced polar codes [364, 357, 365]. Many different variations of polar codes have been proposed in the literature [366, 367, 368, 369]. In particular, polar codes have been examined for 6G-level scenarios, such as ultra-high throughput and extremely reliable low-latency communications [356, 363, 364, 357, 370, 371]. Shaped polar-coded modulation is another active and particularly interesting topic for 6G [360, 63, 361, 362, 372, 373]. The past, present, and future of channel coding were discussed in recent surveys [374, 352, 365]. Further research is needed in these directions to achieve a required level of maturity for practical implementation in the 6G era.
@@
VIII-C3 Non-Orthogonal Multiple Access
@@
  • Literature and Future Directions: Since the early 2010s, NOMA has been widely studied for 5G and beyond [376]. Both domains, power and code, have been considered, with the main focus on the former. As discussed in [376], many NOMA methods were developed and evaluated for 5G NR as a study item. In 3GPP, further study on NOMA was left for beyond 5G scenarios, as its development work was not continued for the work item phase in the 5G NR design process. Since then, a plethora of NOMA schemes have been proposed for promising beyond 5G usage scenarios. Comprehensive surveys on a wide range of future NOMA methods were presented in [378, 381, 382, 383]. In particular, combinations of NOMA with many other emerging 6G technologies have been studied [379].
@@
VIII-C4 Grant-Free Medium Access
@@
    Whereas 4G’s four-step process requires a response from the base station (BS) before sending its payload data, 5G’s two-step access allows devices to transmit data without a particular response [400]. This grant-free random access provides benefits in lowering latency and reducing control signaling overhead, particularly for MTC scenarios with many burst transmissions of small data packets from a massive number of devices [400]. However, the performance of the two-step access is limited by the preamble collision problem when multiple devices send the same preamble to the BS simultaneously, causing a collision of the transmissions. In [402], a survey was provided on the two-step random access procedure in 5G NR, reviewing also the latest literature and potential solutions to the preamble collision problem.
@@
    TABLE VII: Summary of network architectural technologies for 6G
@@
  • Literature and Future Directions: In the literature, there are a handful of survey papers on grant-free access that cover different aspects and review the latest advancements [404, 377, 403, 400, 405, 401]. In [404], compressed sensing techniques were reviewed for grant-free massive IoT connectivity, focusing on the efficient detection of active devices. A comprehensive survey was provided on grant-free NOMA for IoT in [377]. In [403], grant-free URLLC was studied, with a special focus on the potential enhancements provided by massive and cell-free massive MIMO technologies. The paper [400] discussed grant-free random access for MTC. In this context, massive MIMO and NOMA concepts were reviewed, as well as future challenges toward 6G. In [405], the authors explored grant-free access for massive satellite-based IoT in the 6G era. The work [401] discussed grant-free network access based on compressive sensing for 6G massive communications.
@@
VIII-D Network Architectural Technologies for 6G
The design of mobile networks is becoming increasingly complex for each new generation. From the network design perspective, 6G is expected to support 3D network architectures, UDNs, flexible cell deployments, and cell-free operations. A 3D network architecture will be enabled by integrating terrestrial and non-terrestrial access, including space, air, and ground layers. Satellite communication is a key technology for global coverage on the land, in the air, and at the sea. Due to THz communications, 6G will rely on an ultra-dense cell design to provide sufficient coverage for extreme-capacity scenarios. Flexible cell deployments will be facilitated by integrating wireless backhaul and access. In addition to the traditional cell-based design, 6G is expected to be complemented by a cell-free network architecture to provide stable QoS in the coverage area. This section provides a detailed discussion on the aforementioned network architectural technologies, i.e., INTNs, UDNs, IAB, and cell-free massive MIMO. These technologies are summarized in Table VII.
VIII-D1 Integrated Non-Terrestrial and Terrestrial Networks
@@
    Integrating non-terrestrial communications into mobile networks brings many challenges, such as satellite integration at a reasonable cost and QoS, network management in a heterogeneous 3D network architecture, realistic channel modeling, and regulatory aspects. In terms of satellite integration, LEO and mini-satellites are promising candidates for 6G due to their reasonable balance between cost and performance. Satellites in the lower orbits have better propagation features, shorter delays, and cheaper prices than those in the higher orbits. A heterogeneous 3D network architecture with space-air-ground layers makes network and resource management challenging, calling for advanced cooperation mechanisms between different layers. AI/ML is a promising tool for addressing diverse network management issues. Realistic channel models are needed for appropriate performance evaluation in the 6G-specific application scenarios. Regulatory challenges arise when aiming at global coverage. Regulations in different countries and international waters must be considered when designing INTNs. The main challenges of INTNs are further discussed in [407, 411].
  • Literature and Future Directions: In the literature, INTNs have received a considerable amount of attention since the mid-2010s, becoming one of the key topics in 6G research. Numerous survey articles have recently been published [410, 409, 413, 414, 415, 416, 411, 417, 418, 419, 420, 412, 421, 422, 423, 424]. They explored a broad variety of INTN aspects. The work in [410] reviewed 5G NTNs in 3GPP, covering radio access, systems, services, protocols, and IoT. In [409], the authors surveyed the evolution of NTNs from 5G to 6G networks. NTNs were discussed from the perspectives of 5G, mmWave, IoT, multi-access edge computing (MEC), AL/ML, higher layers, field trials, industry progress, and 6G. NTNs were explored toward 6G in [413]. In [414], a survey was conducted on NTNs in the 6G era. The focus was on enabling technologies, open issues, and a case study. 6G SAGIN was studied in [415]. The topics covered include space and mobile networking, key enablers, UAV-as-a-service, design aspects, applications, challenges, and future research avenues.
@@
VIII-D2 Ultra-Dense Networks
@@
  • Literature and Future Directions: In the literature, UDNs have been extensively explored during the past decade [429, 426, 425]. The original driver behind the increased research interest was 5G mmWave communications with dense small-cell deployments [430]. Popular research directions for UDNs have been resource allocation, interference management, user association, energy efficiency, mobility/handover management, and backhauling [429, 427, 426, 425]. Further research is still needed to address these fundamental challenges. Recently, the research focus has shifted to beyond 5G UDNs [431, 428]. Consequently, the variety of studied topics has widened while taking the special characteristics of 6G networks into account. Accordingly, interesting future topics include the exploration of UDNs from the perspective of AI/ML assistance, THz operation, optical wireless, ultra-massive IoT, vehicular networks, ultra-dense IAB, and cell-free network design. In particular, AI/ML is applicable to solve many complex network management problems in the design of UDNs [426, 428, 425]. Further details on UDNs can be found in recent surveys that have reviewed the latest literature, research achievements, open issues, and future directions [427, 426, 428, 425, 432].
@@
VIII-D3 Integrated Access and Backhaul
@@
  • Literature and Future Directions: In recent years, IAB has been extensively studied [433], considering diverse topics from resource and interference management to network topology optimization and end-to-end performance evaluations. A comprehensive overview of the IAB literature was provided in [433]. As the standardization work on NR IAB started in the late 2010s, a special research focus has been directed toward 5G mmWave networks [436]. Further information on NR IAB and the related literature can be found in [223, 434]. Recent IAB topics on 5G evolution and beyond include UAV-aided IAB [437], AI/ML-aided IAB [438], cache-aided IAB [439], RIS-aided IAB [440], mmWave IAB [441], and cell-free IAB [442]. Further discussion of the latest research areas is provided in [433].
    Future IAB research can be divided into two main areas: practical enhancements to 5G NR IAB and IAB solutions beyond 5G. In the former category, essential research directions include enhancements to beamforming, interference management, resource allocation, routing, and topology optimization. For instance, massive MIMO is a potential technology to enhance NR IAB such that the same frequency-time units can be used for access and backhaul links, leading to improved spectral efficiency. Hybrid beamforming is a particularly promising method due to its good balance between performance and implementation complexity. In the latter category, it is important to study the synergy between IAB and other emerging 6G technologies, such as THz IAB, AI/ML-aided IAB, RIS-aided IAB, UAV IAB, cache-enabled IAB, and optical IAB [433]. Since 6G is expected to enter the THz spectrum, a special research focus needs to be directed toward ultra-dense 6G IAB networks, operating at THz frequencies.
@@
VIII-D4 Cell-free Massive MIMO
@@
    In the original concept, all distributed APs are assumed to jointly serve each user in the network area, leading to scalability issues due to the increasing computational complexity and signaling overhead as the number of APs and users grows, making the concept infeasible for practical large-scale implementations. To solve scalability issues, a refined version of the cell-free operation was developed, known as user-centric cell-free massive MIMO [450], in which each user is served by a limited set of nearby APs. Clustering APs for individual users makes the network more scalable, leading to better suitability for large-scale deployments. In [445], the user-centric approach was further enhanced by applying a clustering framework from the network MIMO literature, known as dynamic cooperation clustering. This clustering concept has been proven scalable for any network size [445]. The latest details of cell-free massive MIMO research are covered in [444, 451, 452, 453, 432]. Cell-free massive MIMO is expected to be utilized in 6G to complement the traditional cell-based network architecture in scenarios that require stable QoS in the network coverage area [454, 453, 432].
    TABLE VIII: Summary of network intelligence technologies for 6G
@@
    Since cell-free architecture is a new way to construct a network, resource allocation strategies must be redesigned accordingly. Mobility is also problematic since it leads to rapid chances in clustering. Theoretical channel models are commonly used in the performance evaluation of the proposed cell-free solutions, leading to unrealistic results. Practical channel models based on the dedicated channel measurements are required to achieve more realistic results and more useful performance analyses. Prototyping and field trials are required to shift cell-free massive MIMO from a theoretical concept to a practical network architecture. More detailed discussions of the fundamental challenges of cell-free networks are presented in [444, 451, 452, 453, 432].
  • Literature and Future Directions: In the literature, cell-free massive MIMO has been actively researched since the mid-2010s. The concept has been thoroughly reviewed in recent surveys [455, 456, 454, 444, 451, 452, 457, 458, 459, 453, 432]. A broad range of aspects was discussed, including theoretical foundations, network scalability, dynamic user-centric clustering, synchronization, resource allocation, uplink/downlink operations, AP selection, receive combining, transmit precoding, power optimization, channel hardening, channel estimation, pilot assignment, hardware impairments, mobility issues, mmWave/THz operation, fronthauling, deployment challenges, and future research toward 6G. Although substantial research efforts have been made to date, further work is still needed to overcome the fundamental challenges discussed earlier. Moreover, promising 6G-related future topics include AI/ML for cell-free, cell-free with RISs, cell-free for IoT, cell-free in O-RAN, practical implementations, and experiments [454, 458, 459, 453, 432].
@@
VIII-E Network Intelligence Technologies for 6G
The ultimate goal of AI/ML is to revolutionize the design, operation, and management of mobile networks by making them more intelligent, efficient, flexible, scalable, automated, proactive, economical, ecological, and secure. While this vision is currently far from reality, the first concrete steps toward it have been taken. AI/ML has been adopted in the 5G standards from Release 18 onward (i.e., 5G-Advanced). In 6G, pervasive AI/ML is expected to be a major leap toward the grand vision. This section discusses how AI/ML can be exploited at different levels of the network, i.e., the core, edge, and air interface. These are the cornerstone of 6G network intelligence, as illustrated in Figure 26 and summarized in Table VIII.

### v3

@@ IX-C2 ADVANCED MODULATION AND CODING METHODS
  • Past and Present: As summarized in Figure 20, each mobile generation has adopted a novel or enhanced channel coding concept, i.e., 2G: convolutional coding (data and control), 3G: Turbo coding (data) and convolutional coding (control), 4G: Turbo coding (data) and convolutional coding (control), 5G: low-density parity check (LDPC) coding (data), and polar coding (control) [368]. Currently, LDPC codes are used for data channels and polar codes for control channels in 5G [369]. Turbo coding was the first channel coding method approaching theoretical channel capacity. Linear LDPC coding also belongs to the class of capacity-approaching codes while outperforming Turbo coding, particularly in the higher coding rate region and in terms of the error floor [370]. Polar coding is a linear block-error correction code with low-complexity encoding and decoding designs, and the first channel code that achieves theoretical channel capacity [370]. LDPC and polar codes have been reviewed for 5G NR in [369, 370].
    Figure 20: Evolution of channel coding from 2G to 5G.
@@
    At a high level, there are two main challenges in modulation and coding designs toward 6G, i.e., developing efficient methods for ultra-high throughput and extremely reliable communication links. The former design problem calls for efficient high-order modulation and low coding rate methods with decent error-correcting properties. A straightforward solution is to increase the modulation order beyond 1024-QAM used in 5G, requiring a thorough performance comparison between different variations of QAM. However, increasing the order of QAM may not provide the needed performance gains at higher 6G frequencies due to the inefficiency of power amplifiers [79]. To alleviate this problem, coded modulation with signal shaping is a promising concept [374], especially in combination with polar codes [375, 376, 377, 378]. Polar codes have favorable features to efficiently support lower coding rates in combination with higher-order modulations [375, 379, 380]. Coded modulation with probabilistic shaping is already in use in optical fiber communications [79].
@@
  • Literature and Future Directions: In the 2010s, LDPC and polar codes were studied for 5G [369]. Research beyond 5G is ongoing, with a special emphasis on enhanced polar codes [380, 373, 381]. Many different variations of polar codes have been proposed in the literature [382, 383, 384, 385]. In particular, polar codes have been examined for 6G-level scenarios, such as ultra-high throughput and extremely reliable low-latency communications [372, 379, 380, 373, 386, 387]. Shaped polar-coded modulation has been another active and particularly interesting topic for 6G [376, 79, 377, 378, 388, 389]. The past, present, and future of channel coding were discussed in recent surveys [390, 368, 381]. Further research is needed in these directions to achieve a required level of maturity for practical implementation in the 6G era.
@@
IX-C3 NON-ORTHOGONAL MULTIPLE ACCESS
@@
  • Literature and Future Directions: As discussed earlier, NOMA was not included in the 5G standards but left to be studied further for beyond 5G networks. Since then, a plethora of NOMA schemes have been proposed for promising beyond 5G usage scenarios. Comprehensive surveys on a wide range of future NOMA methods were presented in [394, 397, 398, 399]. In particular, combinations of NOMA with many other emerging 6G technologies were studied in [395].
@@
IX-C4 GRANT-FREE MEDIUM ACCESS
@@
    Whereas the four-step process in 4G requires a response from the base station (BS) before sending its payload data, the two-step access of 5G allows devices to transmit data without a particular response [416]. This grant-free random access provides benefits in lowering latency and reducing control signaling overhead, particularly for MTC scenarios with many burst transmissions of small data packets from a massive number of devices [416]. However, the performance of the two-step access is limited by the preamble collision problem when multiple devices send the same preamble to the BS simultaneously, causing a collision of the transmissions. In [418], a survey was provided on the two-step random access procedure in 5G NR, reviewing also the latest literature and potential solutions to the preamble collision problem.
@@
    TABLE IX: Summary of network architectural technologies for 6G
@@
  • Literature and Future Directions: In the literature, there are a handful of survey papers on grant-free access that cover different aspects and review the latest advancements [393, 419, 416, 420, 417]. A comprehensive survey was provided on grant-free NOMA for IoT in [393]. In [419], grant-free URLLC was studied, with a special focus on the potential enhancements provided by massive and cell-free massive MIMO technologies. The paper [416] discussed grant-free random access for MTC. In this context, massive MIMO and NOMA concepts were reviewed, as well as future challenges toward 6G. In [420], the authors explored grant-free access for massive satellite-based IoT in the 6G era. The work [417] discussed grant-free network access based on compressive sensing for 6G massive communications.
@@
IX-D NETWORK ARCHITECTURAL TECHNOLOGIES FOR 6G
The design of mobile networks is becoming increasingly complex for each new generation. From the network design perspective, 6G is expected to support 3D network architectures, UDNs, flexible cell deployments, and cell-free operations. A 3D network architecture will be enabled by integrating terrestrial and non-terrestrial access, including space, air, and ground layers. Satellite communication is a key technology for global coverage on the land, in the air, and at the sea. Due to THz communications, 6G will rely on an ultra-dense cell design to provide sufficient coverage for extreme-capacity scenarios. Flexible cell deployments will be facilitated by integrating wireless backhaul and access. In addition to the traditional cell-based design, 6G is expected to be complemented by a cell-free network architecture to provide stable QoS in the coverage area. This section provides a detailed discussion on the aforementioned network architectural technologies, i.e., INTNs, UDNs, IAB, and cell-free massive MIMO. These technologies are summarized in Table IX.
IX-D1 INTEGRATED NON-TERRESTRIAL AND TERRESTRIAL NETWORKS
@@
    Integrating non-terrestrial communications into mobile networks brings many challenges, such as satellite integration at a reasonable cost and QoS, network management in a heterogeneous 3D network architecture, realistic channel modeling, and regulatory aspects. In terms of satellite integration, LEO and mini-satellites are promising candidates for 6G due to their reasonable balance between cost and performance. Satellites in the lower orbits have better propagation features, shorter delays, and cheaper prices than those in the higher orbits. A heterogeneous 3D network architecture with space-air-ground layers makes network and resource management challenging, calling for advanced cooperation mechanisms between different layers. AI/ML is a promising tool for addressing diverse network management issues. Realistic channel models are needed for appropriate performance evaluation in the 6G-specific application scenarios. Regulatory challenges arise when aiming at global coverage. Regulations in different countries and international waters must be considered when designing INTNs. The main challenges of INTNs were discussed in [422, 426].
  • Literature and Future Directions: In the literature, INTNs have received a considerable amount of attention since the mid-2010s, becoming one of the key topics in 6G research. Numerous survey articles have recently been published [425, 424, 428, 429, 430, 431, 426, 432, 433, 434, 435, 427, 436, 437, 438, 439]. They explored a broad variety of INTN aspects. The work in [425] reviewed 5G NTNs in 3GPP, covering radio access, systems, services, protocols, and IoT. In [424], the authors surveyed the evolution of NTNs from 5G to 6G networks. NTNs were discussed from the perspectives of 5G, mmWave, IoT, multi-access edge computing (MEC), AL/ML, higher layers, field trials, industry progress, and 6G. NTNs were explored toward 6G in [428]. In [429], a survey was conducted on NTNs in the 6G era. The focus was on enabling technologies, open issues, and a case study. 6G SAGIN was studied in [430]. The main topics included space and mobile networking, key enablers, UAV-as-a-service, design aspects, applications, challenges, and future research avenues.
@@
IX-D2 ULTRA-DENSE NETWORKS
@@
  • Literature and Future Directions: In the literature, UDNs have been extensively explored during the past decade [444, 441, 440]. The original driver behind the increased research interest was 5G mmWave communications with dense small-cell deployments [445]. Popular research directions for UDNs have been resource allocation, interference management, user association, energy efficiency, mobility/handover management, and backhauling [444, 442, 441, 440]. Further research is still needed to address these fundamental challenges. Recently, the research focus has shifted to beyond 5G UDNs [446, 443]. Consequently, the variety of studied topics has widened, while taking the special characteristics of 6G networks into account. Accordingly, interesting future topics include the exploration of UDNs from the perspective of AI/ML assistance, THz operation, optical wireless, ultra-massive IoT, vehicular networks, ultra-dense IAB, and cell-free network design. In particular, AI/ML is applicable to solve many complex network management problems in the design of UDNs [441, 443, 440]. Further details on UDNs can be found in recent surveys that have reviewed the latest literature, research achievements, open issues, and future directions [442, 441, 443, 440, 447].
@@
IX-D3 INTEGRATED ACCESS AND BACKHAUL
@@
  • Literature and Future Directions: In recent years, IAB has been extensively studied [448], considering diverse topics from resource and interference management to network topology optimization and end-to-end performance evaluations. A comprehensive overview of the IAB literature was provided in [448]. As the standardization work on NR IAB started in the late 2010s, a special research focus has been directed toward 5G mmWave networks [451]. Further information on NR IAB and the related literature can be found in [240, 449]. Recent IAB topics on 5G evolution and beyond include UAV-aided IAB [452], AI/ML-aided IAB [453], cache-aided IAB [454], RIS-aided IAB [455], mmWave IAB [456], and cell-free IAB [457]. Further discussion of the latest research areas has been provided in [448].
    Future IAB research can be divided into two main areas: practical enhancements to 5G NR IAB and IAB solutions beyond 5G. In the former category, essential research directions include enhancements to beamforming, interference management, resource allocation, routing, and topology optimization. For instance, massive MIMO is a potential technology to enhance NR IAB such that the same frequency-time units can be used for access and backhaul links, leading to improved spectral efficiency. Hybrid beamforming is a particularly promising method due to its good balance between performance and implementation complexity. In the latter category, it is important to study the synergy between IAB and other emerging 6G technologies, such as THz IAB, AI/ML-aided IAB, RIS-aided IAB, UAV IAB, cache-enabled IAB, and optical IAB [448]. Since 6G is expected to enter the THz spectrum, a special research focus needs to be directed toward ultra-dense 6G IAB networks operating at THz frequencies.
@@
IX-D4 CELL-FREE MASSIVE MIMO
@@
    In the original concept, all distributed APs are assumed to jointly serve each user in the network area, leading to scalability issues due to the increasing computational complexity and signaling overhead as the number of APs and users grows, making the concept infeasible for practical large-scale implementations. To solve scalability issues, a refined version of the cell-free operation was developed, known as user-centric cell-free massive MIMO [465], in which each user is served by a limited set of nearby APs. Clustering APs for individual users makes the network more scalable, leading to better suitability for large-scale deployments. In [460], the user-centric approach was further enhanced by applying a clustering framework from the network MIMO literature, known as dynamic cooperation clustering. This clustering concept has been proven scalable for any network size [460]. The latest details of cell-free massive MIMO research have been discussed in [459, 466, 467, 468, 447]. Cell-free massive MIMO is expected to be utilized in 6G to complement the traditional cell-based network architecture in scenarios that require stable QoS in the network coverage area [469, 468, 447].
    TABLE X: Summary of network intelligence technologies for 6G
@@
    Since cell-free architecture is a new way to construct a network, resource allocation strategies must be redesigned accordingly. Mobility is also problematic since it leads to rapid chances in clustering. Theoretical channel models are commonly used in the performance evaluation of the proposed cell-free solutions, leading to unrealistic results. Practical channel models based on the dedicated channel measurements are required to achieve more realistic results and more useful performance analyses. Prototyping and field trials are required to shift cell-free massive MIMO from a theoretical concept to a practical network architecture. More detailed discussions on the fundamental challenges of cell-free networks can be found in [459, 466, 467, 468, 447].
  • Literature and Future Directions: In the literature, cell-free massive MIMO has been actively researched since the mid-2010s. The concept has been thoroughly reviewed in recent surveys [469, 459, 466, 467, 470, 471, 472, 468, 447]. A broad range of aspects was discussed, including theoretical foundations, network scalability, dynamic user-centric clustering, synchronization, resource allocation, uplink/downlink operations, AP selection, receive combining, transmit precoding, power optimization, channel hardening, channel estimation, pilot assignment, hardware impairments, mobility issues, mmWave/THz operation, fronthauling, deployment challenges, and future research toward 6G. Although substantial research efforts have been made to date, further work is still needed to overcome the fundamental challenges discussed earlier. Moreover, promising 6G-related future topics include AI/ML for cell-free, cell-free with RISs, cell-free for IoT, cell-free in O-RAN, practical implementations, and experiments [469, 471, 472, 468, 447].
@@
IX-E NETWORK INTELLIGENCE TECHNOLOGIES FOR 6G
The ultimate goal of AI/ML is to revolutionize the design, operation, and management of mobile networks by making them more intelligent, efficient, flexible, scalable, automated, proactive, economical, ecological, and secure. While this vision is currently far from reality, the first concrete steps toward it have been taken. AI/ML has been adopted in the 5G standards from Release 18 onward (i.e., 5G-Advanced). In 6G, pervasive AI/ML is expected to be a major leap toward the grand vision. This section discusses how AI/ML can be exploited at different levels of the network, i.e., the core, edge, and air interface. These are the cornerstone of 6G network intelligence, as illustrated in Figure 28 and summarized in Table X.
IX-E1 INTELLIGENT CORE
  • Vision: In the core network, 6G will go beyond the softwarization of 5G by adopting intelligence via pervasive AI/ML. This can be seen as an evolution from cloud to cloud intelligence.
  • Introduction: Mobile networks consist of three main components: core network, RAN, and mobile devices, as illustrated in Figure 29. The core and RAN operate cooperatively between the Internet and mobile devices. The core network is responsible for the overall network operation, management, and security, whereas the RAN handles wireless access to serve mobile devices. For example, one of the key tasks of the core network is to guarantee secure and reliable end-to-end communication between the source and destination nodes. The current core networks are software-based and cloud-native. The next major step in the evolution of core networks is the adoption of intelligence via AI/ML in the 6G era. The intelligent network core is based on the pervasive utilization of AI/ML in order to improve the design, operation, management, maintenance, and security of the network. DL is an attractive technology for the core since it achieves excellent performance when the training datasets are large. Data acquisition is a crucial component of the intelligent core.